# K-theory and localizing invariants of large categories

### v2

@@ Abstract.
We also study the general theory of dualizable categories. In particular, we give an “explicit” proof of Ramzi’s theorem [Ram] on the -presentability of the category of dualizable categories. Among other things, we prove that dualizability is equivalent to “flatness” in the category of presentable stable categories.
@@ 0. Introduction
We call an exact functor of presentable stable -categories strongly continuous if has a right adjoint which is colimit-preserving. We denote by the (non-full) subcategory of dualizable categories and strongly continuous functors. Further, we denote by the subcategory of compactly generated categories.
@@
Another class of examples which we are interested in this paper are categories of sheaves on a locally compact Hausdorff space Let be a presheaf on with values in We denote by the category of sheaves with values in By [Lur18], the category is dualizable. We have the following general result, see 6.10.
@@
In Section 2 we prove a surprising result: a presentable stable category is dualizable if and only if it is flat in More precisely, is flat if the functor preserves fully faithful functors.
In Section 3 we give a criterion of dualizability for categories which are extensions of compactly generated categories. If such an extension is dualizable, then it is in fact compactly generated. As an application, we show that in the extensions of by are classified by functors where is the category of Tate objects, see [Hen17]. As another application, for a commutative noetherian ring we classify the localizing subcategories of which are dualizable. They correspond bijectively to convex subsets of see Theorem 3.10.
In Section 4 we define and study localizing invariants of dualizable categories. In Subsection 4.1 we recall for each regular cardinal the universal localizing invariant of small stable categories, commuting with -filtered colimits. In subsection 4.2 we explain that a localizing invariant of small categories extends uniquely to dualizable categories, using the Calkin construction. In Subsection 4.3 we explain that this canonical extension is in fact the right Kan extension. In Subsection 4.4 we give an example: we compute finitary localizing invariants of the category of sheaves on and on In Subsection 4.5 we do a similar computation for not necessarily finitary localizing invariants, deducing the commutation of K-theory with products (Theorem 4.29).
@@
In Section 6 we study the sheaves on locally compact Hausdorff space with values in a presheaf of dualizable categories. First, in Subsection 6.1 we give a computation of arbitrary localizing invariants for sheaves on finite CW complexes (in the case of constant coefficients). In Subsection 6.2 we explain how to approximate the category of sheaves on an arbitrary compact Hausdorff space by “simpler” categories (Proposition 6.9). Here it is crucial that we use the notion of a -(pre)sheaf from [Lur09]. In Subsection 6.3 we apply the previous results to compute the finitary localizing invariants of categories of sheaves (Theorem 6.10). In Subsection 6.4 we give a simple description of compact objects in the categories of sheaves.
In Appendix A we give the (expected) description of monomorphisms and epimorphisms of presentable and dualizable categories. In Appendix B we show that an image of a homological epimorphism does not have to be a stable subcategory (Proposition B.1). In Appendix C we show that the category is -presentable, and we describe the -compact objects for each uncountable regular (Theorem C.6). The -presentability of is originally due to Ramzi [Ram]. In Appendix D we prove that the category is generated by colimits by a single object – the category of sheaves of spectra on the real line with singular support in In Appendix E we prove an analogue of Adams representability theorem for -compact dualizable categories (Theorem E.11). In Appendix F we explain an analogy between the category of dualizable categories and the opposite category of compact Hausdorff spaces. Finally, in Appendix G we show that -theory of exact -categories commutes with infinite products, using the ideas of Córdova Fedeli [Cor23] and Klemenc [Kle22].
Acknowledgements. I am grateful to Ko Aoki, Alexander Beilinson, Dustin Clausen, Adriano Córdova Fedeli, Vladimir Drinfeld, Boris Feigin, Dennis Gaitsgory, Marc Hoyois, Dmitry Kaledin, David Kazhdan, Maxim Kontsevich, Akhil Mathew, Thomas Nikolaus, Dmitri Orlov, Maxime Ramzi, Victor Saunier, Tomer Schlank, Peter Scholze, Vivek Shende, Vladimir Sosnilo, Georg Tamme, Bertrand Toën, Yakov Varshavsky and Christoph Winges for useful discussions. Part of this work was done while I was a visitor in the Max Planck Institute for Mathematics in Bonn from December 2022 till February 2024, and I am grateful to the institute for their hospitality and support.
@@ 1.1. Infinity-categories, limits and colimits
We will freely use the theory of -categories, functors, limits and colimits as developed in [Lur09, Lur17]. Except for Section 2, we will deal only with -categories. Given an ordinary category we identify it with its nerve which is an -category. In particular, we will consider a poset (partially ordered set) as an -category. Sometimes we will say “category” instead of “-category”, when the meaning is clear from the context.
@@
We refer to [Lur09, Definition 5.3.1.7] for the notion of a -filtered -category. The argument of [AR94, Theorem 1.5, Remark 1.21] shows that for any -filtered -category there exists an -directed poset and a cofinal functor Recall that a poset is called -directed if is nonempty and any -small collection of elements of has an upper bound. Recall that -filtered -categories resp. -directed posets are simply called filtered -categories resp. directed posets.
@@ 1.2. Presentable and accessible stable -categories
An -category is -presentable if is -accessible and cocomplete. Equivalently, this means that where the category has -small colimits. Further, is presentable if it is -presentable for some Following [Lur09], we denote by the -category of all presentable categories and colimit-preserving functors. Note that the category is not locally small.
@@
We denote by the full subcategory of presentable stable categories, and similarly for Note that the category of spectra is an idempotent -algebra in and the subcategory is simply the category of modules over (and similarly for ). In particular, the inclusions and commute with limits and colimits.
@@
. It is sufficient to prove the “moreover” statement. Suppose that and are -presentable and commutes ith -filtered colimuts. Then an object is -compact in whenever We conclude that is -presentable, and the full subcategory is generated by and (as a stable subcategory). ∎
@@
We will need an explicit characterizations of -compact objects in for uncountable
@@
We call an exact functor between presentable stable -categories strongly continuous if it has a right adjoint which is continuous (in other words, if has a twice right adjoint). It is convenient to introduce a non-full subcategory with the same objects and with -morphisms being strongly continuous functors. Note that is locally small. Given presentable stable categories and we denote by the (small) category of strongly continuous functors.
@@
Remark 1.8.
It seems plausible that for any regular cardinal the category is presentable. However, it is not clear how to describe the functor right adjoint to the inclusion.
@@
Given a presentable stable category we will call a full subcategory localizing if is presentable stable and the inclusion functor commutes with colimits. Equivalently, this means that is generated as a cocomplete stable subcategory by a set of objects (or by a single object).
@@ 1.3. Well generated triangulated categories
Recall that for a triangulated category an additive functor is called cohomological if for any exact triangle the sequence is exact. By [Nee01a, Theorem 1.17], a well generated triangulated category satisfies Brown representability for contravariant functors: an additive functor is representable if and only if is cohomological and commutes with products. In particular, an exact functor between well generated triangulated categories has a right adjoint iff commutes with coproducts.
A much more difficult question is the Brown representability for covariant functors, and it is not known if it holds for abstract well generated categories. But if where is presentable stable, then it follows from [Nee09, Theorems 1.11, 1.17 and 1.22] and [Lur09, Proposition A.3.7.6] that an additive functor is corepresentable if and only if is cohomological and commutes with products. In particular, for such and for any well generated triangulated category an exact functor has a left adjoint iff commutes with products.
@@ 1.4. Dualizable categories
Let be another presentable stable category and be a continuous localization, i.e. the right adjoint is fully faithful. Then has a continuous section.
@@
Another description is via the universal problem: we have
@@ 1.5. Examples of compactly assembled categories, stable and non-stable
We refer to Subsection 1.15 for a discussion on general inverse limits in We will study both categories and in great detail in the forthcoming paper [E1]. In particular, we will prove that there is a natural fully faithful strongly continuous functor which induces equivalences on compact objects, and we have
@@
Denote by the category of prenormed -vector spaces, where the morphisms are given by contractive linear maps (i.e. the maps such that ). Then is presentable and compactly assembled. Indeed, the category has colimits, and it is generated by colimits by finite-dimensional normed vector spaces and by the object (-dimensional vector space with zero prenorm). We have
@@ 1.8. Dualizability via compact morphisms
We give a criterion for dualizability of a presentable stable category in terms of the homotopy category At least conceptually, most of this subsection is covered by the results of [Kr00, Kr05], applied to the smashing subcategory (assuming that is dualizable). In particular, the ideal of compact maps corresponds to this smashing subcategory as in [Kr00, Theorem A] and [Kr05, Theorem 1]. However, the criterion for dualizability does not follow formally from the results in loc.cit.
Let be a presentable stable category. We give the following definition.
@@
. Suppose that where is directed, and take some map Dnote by the identity map
@@
. Consider a directed system and a map Since is -compact, we may assume that Indeed, -compactness of implies that for some sequences of elements of the map factors through
@@
We have the usual notion of a tensor product of left and right -modules. In particular, given bimodules we have a bimodule given by Note that the tensor product of quasi-ideals is naturally a quasi-ideal. We say that an bimodule is flat on the right if each right -module is flat,
@@
The inverse map sends an ideal to the quasi-ideal where the map is given by the multiplication.
@@
Let be an idempotent quasi-ideal in and let It is clear that We need to show that as quasi-ideal. Note that if are elements such that or then It follows that the isomorphism factors through hence
Conversely, let be an ideal in such that and is flat on the right. We need to show that is idempotent. It suffices to show that the map
@@
We remark that both Lemma 1.48 and Proposition 1.49 hold in the non-additive resp. non-stable context. Namely, if is any (discrete) small category, then we can consider the (bi)functors with values in the category of sets instead of abelian groups. The tensor product and flatness are defined in the usual way, and the notion of a (quasi-)ideal is the same. The statement of Lemma 1.48 still holds, and the proof is the same.
@@ 1.9. The axiom (AB6)
Let be an -category with infinite products and filtered colimits. We refer to the following condition as the axiom (AB6):
- for any collection of filtered categories and for any functors the map
@@
Suppose that is dualizable. Then the colimit functor has a left adjoint Hence, the functor commutes with infinite products. By Proposition 1.53, this exactly means that (AB6) holds in
Conversely, suppose that holds in Arguing as above, we see that the functor commutes with infinite products. Now, the category is not presentable, so we need the following additional observation to deduce the existence of the left adjoint to
Let be a regular cardinal such that the category is -compactly generated. Denote by the colimit functor. Denoting by the right adjoint to the inclusion, we see that the natural transformation is an equivalence of functors Since both functors and commute with infinite products, we deduce that the functor also commutes with infinite products. Hence, it has a left adjoint Then the composition is left adjoint to hence is dualizable.∎
@@
The same argument shows that a (not necessarily stable) presentable category is compactly assembled if and only if it satisfies (AB6) and the strong version of (AB5): filtered colimits commute with finite limits. If is a complete lattice, then this statement is well known, see [GHKLMS80, Theorem I.2.3].
@@ 1.10. Automatic dualizability and images of functors
We first reduce to the case when consists of a single element. Consider the category Then is dualizable: it is sufficient to prove this for compactly generated categories, and in this case we have By Proposition 1.7, the induced functor is strongly continuous and its image generates
So assume that and put We give two proofs.
First proof. Note that the right adjoint functor is conservative (since the image of generates), and it commutes with products and filtered colimits. Since the axiom (AB6) holds in we deduce that it also holds in By Proposition 1.54, is dualizable.
Second proof. Let be a generating object, and Then for an ind-object we have an equivalence
@@
In the assumptions of Corollary 1.58, if is dualizable, then is also the fiber of the cofiber of in
@@ 1.11. Calkin categories
Fix an uncountable regular cardinal We will define the Calkin and -Calkin categories as follows.
Given a small stable Karoubi complete category we put
It turns out that this construction has a direct analogue for dualizable categories.
Let be a dualizable category. By Corollary 1.22, is -compactly generated. By Corollary 1.23 the image of the functor is contained in the full subcategory Hence, we have a fully faithful strongly continuous functor
@@
in
We have fully faithful functors for We denote by the colimit of
@@
Note that if is compactly generated, then the essential image of the functor is a stable subcategory: it is equivalent to the Verdier quotient However, in general this essential image does not have to be a stable subcategory, see Proposition B.1 below.

### v3

@@ Abstract.
We also study the general theory of dualizable categories. In particular, we give an “explicit” proof of Ramzi’s theorem [Ram24a] on the -presentability of the category of dualizable categories. Among other things, we prove that dualizability is equivalent to “flatness” in the category of presentable stable categories.
@@ 0. Introduction
We call an exact functor of presentable stable -categories strongly continuous if has a right adjoint which is colimit-preserving. We denote by the (non-full) subcategory of dualizable categories and strongly continuous functors. Further, we denote by the full subcategory of compactly generated categories.
@@
Another class of examples which we are interested in this paper are categories of sheaves on a locally compact Hausdorff space Let be a presheaf on with values in We denote by the category of sheaves with values in By [Lur18], the category is dualizable. We have the following general result, see Theorem 6.11.
@@
In Section 2 we prove a surprising result: a presentable stable category is dualizable if and only if it is flat in More precisely, is flat if the functor preserves fully faithful functors. We also give a more general criterion for flatness in for an uncountable regular cardinal
In Section 3 we give a criterion of dualizability for categories which are extensions of compactly generated categories. If such an extension is dualizable, then it is in fact compactly generated. As an application, we recover a result from [CDH+20, Appendix A] stating that in the extensions of by are classified by functors where is the category of Tate objects, see [Hen17]. As another application, for a commutative noetherian ring we classify the localizing subcategories of which are dualizable. They correspond bijectively to convex subsets of see Theorem 3.10.
In Section 4 we define and study localizing invariants of dualizable categories. In Subsection 4.1 we recall for each regular cardinal the universal localizing invariant of small stable categories, commuting with -filtered colimits. In subsection 4.2 we explain that a localizing invariant of small categories extends uniquely to dualizable categories, using the Calkin construction. In Subsection 4.3 we explain that this canonical extension is in fact the right Kan extension. In Subsection 4.4 we give an example: we compute finitary localizing invariants of the category of sheaves on and on In Subsection 4.5 we do a similar computation for not necessarily finitary localizing invariants, deducing the commutation of K-theory with infinite products (Theorem 4.28).
@@
In Section 6 we study the sheaves on locally compact Hausdorff space with values in a presheaf of dualizable categories. First, in Subsection 6.1 we give a computation of arbitrary localizing invariants for sheaves on finite CW complexes (in the case of constant coefficients). In Subsection 6.2 we explain how to approximate the category of sheaves on an arbitrary compact Hausdorff space by “simpler” categories (Proposition 6.10). Here it is crucial that we use the notion of a -(pre)sheaf from [Lur09a]. In Subsection 6.3 we apply the previous results to compute the finitary localizing invariants of categories of sheaves (Theorem 6.11). In Subsection 6.4 we give a simple description of compact objects in the categories of sheaves.
In Appendix A we give the (expected) description of monomorphisms and epimorphisms of presentable and dualizable categories. In Appendix B we show that an image of a homological epimorphism does not have to be a stable subcategory (Proposition B.1). In Appendix C we show that the category is -presentable, and we describe the -compact objects for each uncountable regular cardinal (Theorem C.6). The -presentability of is originally due to Ramzi [Ram24a]. In Appendix D we prove that the category is generated by colimits by the single object – the category of sheaves of spectra on the real line with singular support in In Appendix E we prove two closely related versions of Adams representability theorem for -compact dualizable categories (Theorems E.2 and E.4). In Appendix F we explain an analogy between the category of dualizable categories and the opposite category of compact Hausdorff spaces. Finally, in Appendix G we show that -theory of exact -categories commutes with infinite products, using the results of Córdova Fedeli [Cor23] and Klemenc [Kle22].
Acknowledgements. I am grateful to Ko Aoki, Alexander Beilinson, Dustin Clausen, Adriano Córdova Fedeli, Vladimir Drinfeld, Boris Feigin, Dennis Gaitsgory, Marc Hoyois, Dmitry Kaledin, David Kazhdan, Maxim Kontsevich, Akhil Mathew, Thomas Nikolaus, Dmitri Orlov, Maxime Ramzi, Victor Saunier, Tomer Schlank, Peter Scholze, Vivek Shende, Vladimir Sosnilo, German Stefanich, Georg Tamme, Bertrand Toën, Yakov Varshavsky and Christoph Winges for useful discussions. Part of this work was done while I was a visitor in the Max Planck Institute for Mathematics in Bonn from December 2022 till February 2024, and I am grateful to the institute for their hospitality and support.
@@ 1.1. Infinity-categories, limits and colimits
We will freely use the theory of -categories, functors, limits and colimits as developed in [Lur09a, Lur17]. Except for Section 2, we will deal only with -categories. Given an ordinary category we identify it with its nerve which is an -category. In particular, we will consider a poset (partially ordered set) as an -category. Sometimes we will say “category” instead of “-category”, when the meaning is clear from the context.
@@
We refer to [Lur09a, Definition 5.3.1.7] for the notion of a -filtered -category. By [Lur09a, Proposition 5.3.1.18] for any -filtered -category there exists a -directed poset and a cofinal functor Recall that a poset is called -directed if is nonempty and any -small collection of elements of has an upper bound. Recall that -filtered -categories resp. -directed posets are simply called filtered -categories resp. directed posets.
@@ 1.2. Presentable and accessible stable -categories
An -category is -presentable if is -accessible and cocomplete. Equivalently, this means that where the category has -small colimits. Further, is presentable if it is -presentable for some Following [Lur09a], we denote by the -category of all presentable categories and colimit-preserving functors. Note that the category is not locally small.
@@
We denote by the full subcategory of presentable stable categories, and similarly for The inclusions and commute with limits and colimits.
@@
. It is sufficient to prove the “moreover” statement. Suppose that and are -presentable and commutes with -filtered colimuts. Then an object is -compact in whenever We conclude that is -presentable, and the full subcategory is generated by and (as a stable subcategory). ∎
@@
We will need an explicit characterization of -compact objects in for uncountable
@@
We call an exact functor between presentable stable -categories strongly continuous if it has a right adjoint which is continuous (in other words, if has a twice right adjoint). It is convenient to introduce the non-full subcategory with the same objects and with -morphisms being strongly continuous functors. Note that is locally small. Given presentable stable categories and we denote by the (small) category of strongly continuous functors.
@@
It follows from Proposition 1.72 that the category is not presentable, more precisely, the object is not -compact for any regular cardinal A different example showing this was explained to me by German Stefanich.
@@
Given a presentable stable category we will call a full subcategory localizing if is presentable stable and the inclusion functor commutes with colimits. Equivalently, this means that is generated as a cocomplete stable subcategory by a set of objects (or equivalently by a single object).
@@ 1.3. Well generated triangulated categories
Recall that for a triangulated category an additive functor is called homological if for any exact triangle the sequence is exact. It is convenient to call a functor with the same property cohomological. By [Nee01a, Theorem 1.17], a well generated triangulated category satisfies Brown representability for contravariant functors: an additive functor is representable if and only if is cohomological and commutes with products. In particular, an exact functor between well generated triangulated categories has a right adjoint iff commutes with coproducts.
A much more difficult question is the Brown representability for covariant functors, and it is not known if it holds for abstract well generated categories. But if where is presentable stable, then it follows from [Nee09, Theorems 1.11, 1.17 and 1.22] and [Lur09a, Proposition A.3.7.6] that an additive functor is corepresentable if and only if is homological and commutes with products. In particular, for such and for any well generated triangulated category an exact functor has a left adjoint iff commutes with products.
@@ 1.4. Dualizable categories
Let be another presentable stable category and let be a continuous localization, i.e. the right adjoint is fully faithful. Then has a continuous section.
@@
Another description is via the universal property: we have
@@ 1.5. Examples of compactly assembled categories, stable and non-stable
We refer to Subsection 1.15 for a discussion of general inverse limits in We will study both categories and in great detail in the forthcoming paper [E1]. In particular, we will prove that there is a natural fully faithful strongly continuous functor which induces equivalences on compact objects, and we have
@@
Denote by the category of seminormed -vector spaces, where the morphisms are given by contractive linear maps (i.e. the maps such that ). Then is presentable and compactly assembled. Indeed, the category has colimits, and it is generated by colimits by finite-dimensional normed vector spaces and by the object (-dimensional vector space with zero seminorm). We have
@@ 1.8. Dualizability via compact morphisms
We give a criterion for dualizability of a presentable stable category in terms of the homotopy category At least conceptually, most of this subsection is covered by the results of [Kr00, Kr05], applied to the smashing subcategory (assuming that is dualizable). In particular, the ideal of compact maps corresponds to this smashing subcategory as in [Kr00, Theorem A] and [Kr05, Theorem 1]. However, the criterion for dualizability does not follow formally from the results in loc.cit. Definition 1.38 and a version of Theorem 1.45 below were explained to me by Dustin Clausen. Proposition 1.48 seems to be new.
Let be a presentable stable category.
@@
. Suppose that where is directed, and take some map Denote by the identity map
@@
. Consider a directed system and a map Since is -compact, we may assume that Indeed, -compactness of implies that for some sequence of elements of the map factors through
@@
We have the usual notion of a tensor product of left and right -modules. In particular, given bimodules we have a bimodule given by
Note that the tensor product of quasi-ideals is naturally a quasi-ideal. We say that an bimodule is flat on the right if each right -module is flat,
@@
The inverse map sends an ideal to the quasi-ideal where the map is given by the multiplication (i.e. composition).
@@
Let be an idempotent quasi-ideal in and let It is clear that We need to show that as quasi-ideal. Note that if are elements such that or then It follows that the isomorphism factors through hence Conversely, let be an ideal in such that and is flat on the right. We need to show that is idempotent. It suffices to show that the map
@@
We remark that both Lemma 1.47 and Proposition 1.48 hold in the non-additive resp. non-stable context. Namely, if is any (discrete) small category, then we can consider the (bi)functors with values in the category of sets instead of abelian groups. The tensor product and flatness are defined in the usual way, and the notion of a (quasi-)ideal is the same. The statement of Lemma 1.47 still holds, and the proof is the same. If is a compactly assembled accessible -category (not necessarily cocomplete), then the analogue of Proposition 1.48 still holds, and the proof is the same. If is an ordinary category, then is the quasi-ideal of wavy arrows in the terminology of [JJ82, Section 2].
@@ 1.9. The axiom (AB6)
Let be an -category with infinite products and filtered colimits. We refer to the following condition as the axiom (AB6): - for any collection of filtered categories and for any functors the map
@@
Suppose that is dualizable. Then the colimit functor has a left adjoint Hence, the functor commutes with infinite products. By Proposition 1.52, this exactly means that (AB6) holds in Conversely, suppose that holds in Arguing as above, we see that the functor commutes with infinite products. Now, the category is not presentable, so we need the following additional observation to deduce the existence of the left adjoint to Let be a regular cardinal such that the category is -compactly generated. Denote by the colimit functor. Denoting by the right adjoint to the inclusion, we see that the natural transformation is an equivalence of functors Since both functors and commute with infinite products, we deduce that the functor also commutes with infinite products. Hence, it has a left adjoint Then the composition is left adjoint to hence is dualizable.∎
@@
The same argument shows that a (not necessarily stable) presentable category is compactly assembled if and only if it satisfies (AB6) and the strong version of (AB5): filtered colimits commute with finite limits. If is a complete lattice, then this statement is well known, see [GHKLMS80, Theorem I.2.3]. Below we will see that the category satisfies (AB6) (Proposition 1.106) but the strong (AB5) does not hold in (Corollary 1.86).
@@ 1.10. Automatic dualizability and images of functors
We first reduce to the case when consists of a single element. Consider the category Then is dualizable: it is sufficient to prove this for compactly generated categories, and in this case we have By Proposition 1.7, the induced functor is strongly continuous and its image generates So assume that and put We give two proofs. First proof. Note that the right adjoint functor is conservative (since the image of generates), and it commutes with products and filtered colimits. Since the axiom (AB6) holds in we deduce that it also holds in By Proposition 1.53, is dualizable. Second proof. Let be a generating object, and Then for an ind-object we have an equivalence
@@
Under the assumptions of Corollary 1.57, if is dualizable, then is also the fiber of the cofiber of in
@@ 1.11. Calkin categories
Fix an uncountable regular cardinal We will define the Calkin and -Calkin categories as follows. Given a small stable Karoubi complete category we put
It turns out that this construction has a direct analogue for dualizable categories. Let be a dualizable category. By Corollary 1.21, is -compactly generated. By Corollary 1.22 the image of the functor is contained in the full subcategory Hence, we have a fully faithful strongly continuous functor
@@
in We have fully faithful functors for We denote by the colimit of
@@
Note that if is compactly generated, then the essential image of the functor is a stable subcategory: it is equivalent to the Verdier quotient However, in general this essential image does not have to be a stable subcategory even for see Proposition B.1 below.